OpenStudy (anonymous):

Medal For $115, a pharmacist purchases 5 L of cough syrup and repackages it in 125-milliliter bottles. Each bottle costs the pharmacist $0.25. Each bottle of cough syrup is sold for $23.69. Find the profit on the 5 L of cough syrup. (Round your answer to the nearest cent.) $ ?

OpenStudy (anonymous):

what grade is this stuff?

OpenStudy (anonymous):

Grade 12

OpenStudy (anonymous):

oh well im 10th so

OpenStudy (anonymous):

oh ok

OpenStudy (princeharryyy):

total bottles he can make = 5000ml/125ml = 40bottles cost of 40 bottles = 40*$0.25 = $10 bottles sold at = 40*$23.69 = $947.6 profit($) = sp -cp => $947.6-$10 = $937.6

OpenStudy (princeharryyy):

convert $938 in cents that will give u ur answer @harz360

OpenStudy (anonymous):

$938.60

OpenStudy (princeharryyy):

wats dis

OpenStudy (anonymous):

how would you convert to cents convert what to cents

OpenStudy (princeharryyy):

wait...)) i misunderstood the question a bit.. let me solve it again @harz360

OpenStudy (anonymous):

okay

OpenStudy (princeharryyy):

total bottles he can make = 5000ml/125ml = 40bottles cost of 40 bottles = 40*$0.25 = $10 as bottles required is 40 and the cost of those bottles is $10, total cost price for pharmacist = cost of syrup + cost of bottles = 115+10 = $125 40 bottles sold at = 40*$23.69 = $947.6 profit($) = sp -cp => $947.6-$125 = $822.6 would be the answer... :) n SORRY for the earlier answer... @harz360

OpenStudy (anonymous):

so the answer is $822.60 and no its alright

OpenStudy (princeharryyy):

yes...! ohk! @harz360

OpenStudy (anonymous):

thank you :)

OpenStudy (princeharryyy):

but i am not getting have they asked the answer in dollars or cents... if they have asked in cents just multiply it by 100 as 1 dollar equals 100 cents i think.... u must be knowing that.. @harz360